In the Eruopean Land snail, multipe alleles at a single locus determine shell color. The allele for brown (SB) is dominant to the allele for pink (SP) and to the allele for yellow (SY). PInk is recessive to brown, but dominant to yellow.

a. In a population of 1000 snails, 160 are yellow. Assuming the populations is Hardy-Weinberg Equilibrium, calculate th frequency of the yellow allele,

r = f(SY) = ?

b. Given that p = f(SB) = 1. Calculate the frequency of each of the following. Show work please

f (SBSB) = ?

f (SBSP) =?

f (SBSY) = ?

f (SPSP) = ?

f(SPSY) = ?

f (SYSY) = ?

c. Give the phenotypic frequencies and show all work please

f(brown) = ?

f(pink) = ?

f(yellow) = ?
Answer #1

a. In a population of 1000 snails, 160 are yellow. Assuming the populations is Hardy-Weinberg Equilibrium, calculate
the frequency of the yellow allele (recessive) is r = f(SY) = 160/1000 = 0.16 (homozygous recessive)
therefore, the frequency of the allele for yellow shells q = 0.4, which means p=0.6.

b. Frequency of SBSB = 0.6*0.6 = 0.36
Frequency of SBSP = 0.6*0.4 = 0.24
Frequency of SBSY = 2*0.6*0.4 = 0.4